Needhelp123456:

Simplify the following expression. cot^2 x sec x-cos x

jhonyy9:

like a first step you need to know that sec x = 1/cos x cot x = cos x / sin x so cot^2 x = cos^2 x / sin^2 x using these above wrote you can rewriting the given equation in this form cot^2 x sec x - cos x = = (cos^2 x / sin^2 x ) * (1/cos x ) - cos x = = (cos x /sin^2 x ) - cos x = = [cos x /(1-cos^2 x)] - cos x hope helped